OpenStudy (anonymous):

Take a break from integrals: Find the remainder when \(\sum\limits_{n=1}^{2015}n!\) is divided by \(15\).

OpenStudy (mathmath333):

as \(\large\tt \begin{align} \color{black}{5!+6!+7!+............2015!\hspace{1.5em}\\ }\end{align}\) all are divisible by \(\large\tt \begin{align} \color{black}{15\hspace{1.5em}\\}\end{align}\) so \(\large\tt \begin{align} \color{black}{\dfrac{1!+2!+3!+4!+5!.......2015!}{15}\hspace{1.5em}\\~\\ \implies =\dfrac{1!+2!+3!+4!}{15}\hspace{1.5em}\hspace{1.5em}\\~\\ \implies =\dfrac{33}{15}\hspace{1.5em}\hspace{1.5em}\\~\\~\\ \Large \implies \equiv3\hspace{1.5em}\hspace{1.5em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

That's correct!